# FinPos: A Position-Aware Trading Agent System for Real Financial Markets

### v1

@@ Abstract.
The exceptional potential of large language models (LLMs) in handling text information has garnered significant attention in the field of financial trading. However, current trading agents primarily focus on single-step trading tasks and lack awareness of continuous position management. Therefore, we propose a position-aware trading task designed to simulate a more realistic market. To address this task, we develop a trading agent system, FinPos, optimized for position management. FinPos is able to interpret various types of market information from a professional perspective, providing a reliable basis for positioning decisions. To mitigate the substantial market risks arising from position fluctuations, FinPos employs dual decision agents. Furthermore, the continuous nature of position management necessitates our adoption of multi-timescale rewards, which in turn empowers FinPos to effectively balance short-term fluctuations against long-term trends. Extensive experiments demonstrate that FinPos surpasses state-of-the-art trading agents in the position-aware trading task, which closely mirrors real market conditions. More importantly, our findings reveal that LLM-centered agent systems exhibit a vast, largely unexplored potential in long-term market decision-making.
Key words and phrases:
@@
Trading Agent, Position Aware, Real Markets.
Corresponding author.
1. Introduction
@@
Figure 1. With the introduction of position awareness, the agent must not only predict current market trends but also manage the remaining positions in the account. Agents developed for tasks without position awareness are inadequate for addressing the new challenges posed by position-aware trading tasks.
As communication and information technologies evolve, the financial markets are seeing an unprecedented surge in critical, market-moving information Revsine et al. (2021). Due to inherent cognitive limitations, human traders struggle to process the sheer volume of financial data efficiently Black (1986), which poses a serious challenge for real-time market responses. As a result, automated trading systems Huang et al. (2019); Tran et al. (2023); Adediran et al. (2025), which leverage sophisticated models, are becoming increasingly vital within the financial industry.
In the past decade, deep reinforcement learning (DRL) agents Fischer (2018); Shakya et al. (2023) gained considerable attention due to their ability to handle large financial datasets. These agents utilize deep neural networks Millea (2021) to extract meaningful features, enabling them to approximate complex market dynamics. While DRL agents can learn strategies to maximize investment returns, they still struggle to incorporate unstructured textual information—such as news articles Lillo et al. (2015) and financial reports Oyewole et al. (2024). Furthermore, their limited interpretability Balhara et al. (2022) makes it difficult for human traders to trust or collaborate with them effectively.
@@
Recently, large language models (LLMs) Hurst et al. (2024); Liu et al. (2024); Yang et al. (2025) have demonstrated remarkable potential in handling text-modal information and solving complex tasks, such as reasoning and decision-making Zhai et al. (2025). This suggests that LLM-centered agent systems have the potential to transcend the inherent capability boundaries of traditional DRL trading models. Specialized LLM agents designed for financial trading, like FinGPT Yang et al. (2023), FINMEM Yu et al. (2024a), and FinAgent Zhang et al. (2024b), have already achieved groundbreaking results. However, as shown in the left panel of Fig. 1, current LLM agents primarily tackle single-step trading tasks, making them more akin to isolated simulations than realistic financial decision-making systems. In real markets, balancing risk and return requires a sophisticated trading agent to shift its focus from short-term market fluctuations to long-term position management. Therefore, we propose a position-aware trading task, which more closely aligns with real-world trading markets than previous setups.
@@
Existing agent architectures, as shown in Fig. 1, struggle to sustain high yield rates when tasked with position management. This is mainly because they lack three essential capabilities: (1) Position-based risk management: When an agent is responsible for determining investment scale, the associated risks of its decisions are considerably greater than those encountered in daily trend forecasting tasks. (2) Long-term planning capability: Trading tasks that are re-initialized daily are prone to developing agents with a myopic trading perspective. Effective position management, however, necessitates a more long-term investment strategy. (3) In-depth market analysis ability: Short-term trends are often identifiable through market sentiment, yet a much deeper analysis is essential for understanding long-term growth potential. To address these challenges, we propose FinPos, a trading agent capable of managing portfolio positions within real financial markets.
To address position-based risk management, FinPos adopts a multi-agent decision-making system that separates the direction decision agent (buy/sell/hold) from the scale decision agent (transaction volume). This decoupling makes risk management more flexible and controllable. The investment direction decision agent is exclusively tasked with determining whether to buy, sell, or hold on each trading day. Meanwhile, the investment scale decision agent dictates the specific transaction volume to buy or sell, subject to an upper limit on single transaction sizes to ensure capital safety. Additionally, we incorporate risk-awareness and position-management awareness into the investment scale decision agent via prompts, keeping investment risk within controllable bounds. More importantly, as financial market risk originates not only from firm-specific events but also from macroeconomic developments (e.g., tariff regimes), FinPos incorporates macroeconomic news monitoring to proactively hedge against systemic risks.
Long-term planning capability is an essential requirement for a position-aware trading agent. Current financial LLM agents are limited by daily trading tasks, leading them to exclusively focus on the next day’s price movements. This limitation results in agents adopting narrow, day-to-day trading strategies. In real markets, agents requires not only the ability to anticipate short-term fluctuations but also a long-term perspective on assets. To address this, we design a comprehensive short-term and long-term reward framework for FinPos. First, we employ a window-based design that evaluates today’s decisions against outcomes over multiple days, ensuring that short-term actions are judged in a longer context. Second, we embed a multi-timescale reward into the reflection process, so that after each action the agent re-examines its choice and measures consistency with both immediate and cumulative returns. Together, these components allow FinPos to move beyond “one-day-at-a-time” strategies and acquire genuine long-term planning ability during training.
@@
By observing the reasoning chain of trading agents, we observe that current trading agents often analyze market information superficially and lack causal reasoning abilities. For instance, when trading Tesla stock, the agent may overestimate the relevance of news such as ”Musk invests in company A” while dismissing significant news like ”major changes in U.S. tariff policy” as irrelevant to trading decisions. This issue arises because the information analysis agent lacks fundamental common sense about the specific forms of information. To enhance in-depth market analysis ability, we employ a strategy of independent analysis by multiple agents for various types of market information (e.g., financial reports, news) and inject fundamental domain-specific knowledge and focused guidance on key issues into each analyst. With this mechanism, FinPos achieves deeper insights into the market.
@@
2. Related Works
2.1. LLM-Based AI Agents
As researchers increasingly recognize the extraordinary capabilities of large language models (LLMs) in tasks such as summarization Zhang et al. (2024a), reasoning Ke et al. (2025), and decision-making Eigner and Händler (2024), AI agents built around LLMs are gaining significant attention. These LLM agents operate as closed-loop systems, autonomously gathering information, storing memories, reasoning, making decisions, and observing environmental feedback. The advancement of LLMs is propelling AI agents into diverse fields, from open-ended artistic creation to precise medical applications. For instance, MusicAgent Yu et al. (2023) combines various music-related tools and autonomous workflows to assist artists in music creation. MedAgents Wang et al. (2025) use a multi-disciplinary collaboration framework to improve the efficiency of medical diagnosis. Programmer Agent automates the development of software projects. ResearchAgent Li et al. (2024b) helps researchers generate innovative ideas from extensive literature reviews. Personal assistant agents Li et al. (2024a) engage closely with users, providing solutions for tasks and emotional support. These applications illustrate the vast potential of LLM agents in multiple facets of human society. We aim to bring this potential to the financial markets, making LLM agents key players in investment activities.
2.2. Financial LLM-Based Agents
The evolution of financial agent architectures has exhibited significant progress through successive innovations and expansions in cognitive modeling and system integration. FINMEM Yu et al. (2024a) was the first to establish a fundamental trading agent architecture centered around analysis, memory, and decision modules. Building upon this foundation, FinAgent Zhang et al. (2024b) and FinVision Fatemi and Hu (2024) integrate multimodal data—encompassing numerical, textual, and visual information—to analyze both dynamic financial market trends and historical trading patterns. FINCON Yu et al. (2024b) and TradingAgents Xiao et al. (2024) introduced the concept of investment group architectures into agent systems, employing multi-agent collaboration mechanisms to enhance decision-making reliability. Fundamentally, these agents primarily address a simplified financial game focused on predicting next-day price movement. In real markets, however, agents must not only needs to determine the investment direction, but also decide when to adjust exposure and how much to allocate, considering a comprehensive historical investment context. Therefore, this paper adopts a position-aware trading task that more closely approximates real market environments, enabling a more authentic validation of trading agents’ capabilities.
3. Task Definition
3.1. Single-Step Trading Task
@@
This setup is widely adopted in studies such as FINMEM Yu et al. (2024a) due to its simplification of trading processes, which stabilizes the training environment and allows more straightforward model evaluation. However, this simplification comes with significant limitations: it completely ignores the continuity of holding positions and the gains and losses arising from holding positions in real trading, thus cannot capture the associated compound returns or the risk exposure carried across periods. Without modeling positions, the agent has no basis for risk control or for detecting market trends, both of which are central to real trading.These omissions weaken the agent’s learning and generalization in essential skills such as position management and risk control.More critically, by reducing trading to short-term, step-by-step predictions, such a setup underutilizes the natural strengths of LLMs in reasoning over longer horizons. Consequently, advancing financial trading agents towards practical application requires a task that more closely mirrors real market conditions.
3.2. Position-Aware Trading Task
@@
$R=\sum_{t=1}^{N}r_{t},\quad r_{t}=\text{position}_{t}\times\log\frac{\text{price}[t+1]}{\text{price}[t]}$ (2)
Where $\text{position}_{t}$ represents the agent’s current position state, dynamically updated based on past trading decisions. This setup not only more accurately models the continuity and trend dynamics of holding positions in trading but also allows the agent to naturally incorporate key capabilities in strategy learning, such as position control, risk exposure management, and entry-exit timing. The Position-Aware Trading Task provides the agent with a more challenging and practically meaningful learning objective, offering stronger potential for strategy generalization and real-world applicability.
@@ 4. Architecture of FinPos
The architecture of FinPos (Fig. 2) consists of three core modules: the market signal processing and analysis module, the trading decision module, and the multi-timescale reward reflection module. The signal processing module introduces a two-level agent hierarchy—filtering and analysis—to process raw data, with the results stored in a memory layer. The decision module advances beyond prior research that only predicts direction by introducing a position-aware mechanism. A direction decision agent determines the trading orientation, while a quantity decision agent dynamically adjusts position size in response to market risk exposure. The reward module adopts a multi-timescale reward feedback mechanism to enhance the long-term behavioral stability of the model and its responsiveness to trading opportunities. The details of these three modules are presented in the subsequent sections, while all agent prompts are provided in the Appendix A.
@@
Figure 2. Architectural Details of FinPos: Initially, multiple analysis agents leverage domain knowledge to gather diverse information from the environment, subsequently storing it in the memory module. The memory module utilizes a memory allocator to distribute the acquired information across memory layers of varying depths. Subsequently, the most pertinent information for the current decision is placed into working memory, where dual decision agents output trading actions. Finally, multi-timescale rewards guide reflection, storing experiential knowledge into deeper memory layers.
4.1. Market Signal Processing and Analysis
@@
To ensure task clarity and enable LLMs to perform effectively in complex market environments, FinPos draws inspiration from institutional investment workflows. It assigns specialized agents to distinct information domains—such as macroeconomics, industry trends, and corporate fundamentals—forming a division of labor similar to that in hedge funds or private equity firms. For each domain, we further establish a two-tier structure: Signal Processing Agents and Analysis Agents.
@@
Signal Processing Agents: These agents are responsible for preprocessing and filtering high-noise, heterogeneous market data streams. We observe that sentiment analysis algorithms often produce extreme or even misleading judgments; therefore, leveraging LLMs’ superior language understanding for data filtering proves more reliable. Through domain-specific prompts and heuristic rules, the agents clean, compress, and prioritize information by relevance and importance. For example, large volumes of low-value or weakly related news are downweighted or discarded, while impactful macroeconomic policies or firm-level events are highlighted and passed on to subsequent modules.
@@
Analysis Agents: These agents operate on the filtered data. Each is designed to address a specific information domain and is enhanced with relevant financial knowledge. We observe that large language models often exhibit financial hallucinations in market scenarios, making it difficult for them to accurately capture the key factors influencing stock prices. They tend to rely solely on explicit keywords in news while overlooking the underlying market logic. For example, an LLM may fail to recognize that news about the S&P 500 is strongly correlated with the performance of leading U.S. stocks. To mitigate this issue, we explicitly inject financial knowledge into the prompts, thereby strengthening their causal awareness. As financial reasoning is gradually integrated into the prompts, the agents evolve from mechanical summarizers into analysts with genuine financial reasoning ability, capable of generating deeper insights and producing more accurate interpretations of market dynamics.
@@
4.2. Dual Trading Decision
In real-world trading, an account’s current holdings directly reflect its risk exposure. Professional traders typically adjust their buy–sell decisions dynamically based on existing positions to maintain a relative balance between risk and return. Following this principle, FinPos adopts a dual-agent framework that embeds position awareness into both the Direction Decision Agent and the Quantity and Risk Decision Agent. By integrating multi-source information retrieved from the memory layer, the dual-agent framework enables a decision-making mechanism that more closely mirrors practical trading logic.
4.2.1. The Direction Decision Agent
Incorporating position awareness requires integrating longer-term information. Unlike conventional direction agents that only focus on individual trades, our agent combines structured memory (news, financial reports, and past actions) with the current portfolio state and explicitly specifies the strategic intent of each action—whether it represents a long-term position-building move or a short-term tactical adjustment to exploit local trends. This explicit articulation not only enhances interpretability for human supervisors but also provides downstream quantity agents with a clear strategic context. To prevent the agent from over-relying on short-term fluctuations, we design prompts that encourage it to factor in longer horizons, thereby gradually cultivating its long-term planning capability.
4.2.2. Quantity and Risk Decision Agent
Most existing trading agents assume fixed-unit trades, overlooking the dynamic adjustment of position size and risk exposure. To overcome this limitation, FinPos introduces the Quantity and Risk Decision Agent, which determines the specific trade size after a directional decision has been made. This agent considers the account’s position state, structured memory, and risk references (e.g. Conditional Value at Risk, CVaR Rockafellar et al. (2000); see Appendix B.1) to exercise more precise control over capital allocation and risk levels. In addition, to prevent excessive exposure, we cap the size of each trade to the maximum suggested by the 95% CVaR. Such a mechanism improves adaptability under market volatility and supports balancing short-term opportunities with long-term robustness.
4.3. Multi-Timescale Reward Design

### v2

@@ Abstract
The exceptional potential of large language models (LLMs) in handling text information has garnered significant attention in the field of financial trading. However, most existing trading agents operate under intraday, independent unit-based trading tasks, where decisions are made as isolated directional actions, and thus lack awareness of continuous position management. Therefore, we propose a position-aware trading task designed to simulate a more realistic market. To address this task, we propose FinPos, a position-aware trading agent system designed to explicitly model and manage continuous positions. FinPos enhances position awareness through three key mechanisms: (1) professional-level interpretation of heterogeneous market information; (2) a dual-agent decision structure that separates directional reasoning from risk-aware position adjustment; and (3) multi-timescale reward signals, allowing the agent to internalize position awareness through experiential feedback rather than static instructions alone. Extensive experiments demonstrate that FinPos surpasses state-of-the-art trading agents in the position-aware trading task, which closely mirrors real market conditions. More importantly, our findings reveal that LLM-centered agent systems exhibit a vast, largely unexplored potential in long-term market decision-making.
@@
FinPos: A Position-Aware Trading Agent System for Real Financial Markets
Bijia Liu Dongbei University of Finance and Economics qq1403492787@gmail.com          Ronghao Dang Alibaba DAMO Academy dangronghao.drh@alibaba-inc.com
1 Introduction
@@
Figure 1: With the introduction of position awareness, the agent must not only predict current market trends but also manage the remaining positions in the account. Agents developed for tasks without position awareness are inadequate for addressing the new challenges posed by position-aware trading tasks.
An explosion of market information, facilitated by evolving communication technologies, highlights the growing significance of automated trading systems. In the past decade, deep reinforcement learning (DRL) agents Fischer (2018); Shakya et al. (2023) gained considerable attention due to their ability to handle large financial datasets. However, their inherent difficulty in integrating unstructured text information Benk (2022) limits their applicability primarily to technical analysis and compromises their interpretability Balhara et al. (2022).
@@
Recently, large language models (LLMs) Hurst et al. (2024); Liu et al. (2024) have demonstrated remarkable potential in handling text-modal information and solving complex tasks, such as reasoning and decision-making Zhai et al. (2025). This suggests that LLM-centered agent systems have the potential to transcend the inherent capability boundaries of traditional DRL trading models. Specialized LLM agents designed for financial trading, like TradingAgents Xiao et al. (2024), FINMEM Yu et al. (2024a), and FinAgent Zhang et al. (2024), have already achieved groundbreaking results. However, as shown in the left panel of Fig. 1, most existing LLM-based trading agents are formulated under single-step trading tasks: each day is treated as an isolated decision episode, with discrete actions and fixed unit sizes. Although some studies mention "position" in prompts, the single-step trading tasks fundamentally undermines continuous position management. In contrast, real-world trading is inherently multi-step and stateful: a trader’s current holdings directly shape future risk, opportunity, and decision logic. To bridge this gap, we propose a position-aware trading task, which more closely aligns with real-world trading markets than previous setups.
@@
As shown in the right panel of Fig. 1, existing agent architectures struggle to sustain high yield rates when tasked with position management. This is mainly because they lack three essential capabilities: (1) Explicit position representation and exposure control: Without a continuous position state, existing agents cannot perform risk-aware, dynamic exposure control. (2) Long-term planning capability: Trading tasks that are re-initialized daily are prone to developing agents with a myopic trading perspective. (3) In-depth market analysis ability: Short-term trends are often identifiable through market sentiment, yet a much deeper analysis is essential for understanding long-term growth potential.
To address these challenges, we propose FinPos, a position-aware trading agent designed to manage portfolio positions in realistic financial markets. FinPos integrates a dual-agent decision architecture with multi-timescale reward mechanism, where a Direction Decision Agent selects discrete actions (buy/sell/hold) and a Quantity and Risk Decision Agent determines transaction volumes.
A defining challenge in long-term planning capability is that the risk and return consequences of an exposure decision are inherently delayed and cumulative—they unfold over multiple trading days rather than within a single step. To explicitly encode this temporal dependency into the agent’s reasoning, FinPos adopts a multi-timescale reward market signal during training, which jointly considers daily immediate profit-and-loss (PnL) and medium- to long-term cumulative returns. This guides the Direction Agent to make non-myopic action choices that balance short-term responsiveness with strategic consistency.
To address explicit position representation and exposure control, the multi-timescale reward framework also feeds long-horizon performance signals into the reflection process of the Quantity and Risk Decision Agent. This enables the agent to internalize the concept of continuous position and understand how its sizing decisions impact portfolio dynamics. Consequently, the Quantity and Risk Decision Agent dynamically determines the specific transaction volume to buy or sell—subject to a CVaR-based upper bound on single-transaction exposure to ensure capital safety.
@@
By observing the reasoning chain of trading agents, we observe that current trading agents often analyze market information superficially and lack causal reasoning abilities. For instance, when trading Tesla stock, the agent may overestimate the relevance of news such as "Musk invests in company A" while dismissing significant news like "major changes in U.S. tariff policy" as irrelevant to trading decisions. This issue arises because the information analysis agent lacks fundamental common sense about the specific forms of information. To enhance in-depth market analysis ability, we build upon the multi-agent information processing paradigm, go further by by injecting domain-specific financial knowledge and causal principles through tailored prompting. This enables FinPos to interpret market signals not just by sentiment, but by fundamental relevance and risk implications.
@@
2 Related Works
2.1 LLM-Based AI Agents
The evolution of LLMs capabilities has led to a growing interest in LLM-based agents. Recently, the LLM agents have expanded into various domains, including music Yu et al. (2023), healthcare Wang et al. (2025), and research Li et al. (2024). These applications illustrate the vast potential of LLM agents in multiple facets of human society. We aim to bring this potential to the financial markets, making LLM agents key players in investment activities.
2.2 Financial LLM-Based Agents
The evolution of financial agent architectures has exhibited significant progress through successive innovations and expansions in cognitive modeling and system integration. FINMEM Yu et al. (2024a) was the first to establish a fundamental trading agent architecture centered around analysis, memory, and decision modules. Building upon this foundation, FinAgent Zhang et al. (2024) integrate multimodal input to enrich their information sources. FINCON Yu et al. (2024b) and TradingAgents Xiao et al. (2024) introduced the concept of investment group architectures into agent systems to enhance decision-making reliability. Fundamentally, these agents primarily address a simplified financial game focused on predicting next-day price movement. In real markets, however, agents must not only needs to determine the investment direction, but also decide when to adjust exposure and how much to allocate. Therefore, this paper adopts a position-aware trading task that more closely approximates real market environments.
3 Task Definition
3.1 Single-Step Trading Task
@@
This setup is widely adopted in studies such as FINMEM Yu et al. (2024a) due to its simplification of trading processes, which stabilizes the training environment and allows more straightforward model evaluation. However, this simplification comes with significant limitations: it completely ignores the continuity of holding positions and the gains and losses arising from holding positions in real trading, thus cannot capture the associated compound returns or the risk exposure carried across periods. Without modeling positions, the agent has no basis for risk control or for detecting market trends, both of which are central to real trading. More critically, by reducing trading to short-term, step-by-step predictions, such a setup underutilizes the natural strengths of LLMs in reasoning over longer horizons. Consequently, advancing financial trading agents towards practical application requires a task more closely mirrors real market conditions.
3.2 Position-Aware Trading Task
@@
$R=\sum_{t=1}^{N}r_{t},\quad r_{t}=\text{pos}_{t}\times\log\frac{\text{price}[t+1]}{\text{price}[t]}$ (2)
Where $\text{pos}_{t}$ represents the agent’s current position state, dynamically updated based on past trading decisions. This setup not only more accurately models the continuity and trend dynamics of holding positions in trading but also allows the agent to naturally incorporate key capabilities in strategy learning, such as position control, risk exposure management, and entry-exit timing. The Position-Aware Trading Task provides the agent with a more challenging and practically meaningful learning objective, offering stronger potential for strategy generalization and real-world applicability.
@@ 4 Architecture of FinPos
The architecture of FinPos (Fig. 2) consists of three core modules. The market signal processing module adopts a two-level agent hierarchy, consisting of filtering agents and analyst agents, to distill raw market data into structured, decision-relevant signals. The trading decision module employs a dual decision framework that explicitly decouples quantity decision-making from directional determination. The multi-timescale reward module adopts a multi-timescale reward feedback mechanism to experientially internalize the long-term risk and return implications of position changes. The details of these three modules are presented in the subsequent sections, while all agent prompts are provided in the Appendix  A.
@@
Figure 2: Architectural Details of FinPos: Initially, multiple analysis agents leverage domain knowledge to gather diverse information from the environment, subsequently storing it in the memory module. The memory module utilizes a memory allocator to distribute the acquired information across memory layers of varying depths. Subsequently, the most pertinent information for the current decision is placed into working memory, where dual decision agents generate trading actions, while multi-timescale reward signals support reflective updates that consolidate experiential knowledge into deeper memory layers.
4.1 Market Signal Processing and Analysis
@@
FinPos draws inspiration from institutional investment workflows. It assigns specialized agents to distinct information domains forming a division of labor similar to that in private equity firms. For each domain, we further establish a two-tier structure: Signal Processing Agents and Analysis Agents.
@@
Signal Processing Agents: These agents are responsible for preprocessing and filtering high-noise, heterogeneous market data streams. Through domain-specific prompts and heuristic rules, the agents clean, compress, and prioritize information by relevance and importance. For example, large volumes of low-value or weakly related news are downweighted or discarded, while impactful macroeconomic policies or firm-level events are highlighted and passed on to subsequent modules.
@@
Analysis Agents: These agents analysis the filtered data. We observe that large language models often exhibit financial hallucinations in market scenarios, making it difficult for them to accurately capture the key factors influencing stock prices. They tend to rely solely on explicit keywords in news while overlooking the underlying market logic. For example, an LLM may fail to recognize that news about the S&P 500 is strongly correlated with the performance of leading U.S. stocks. To mitigate this issue, we explicitly inject financial knowledge into the prompts, thereby strengthening their causal awareness. As financial reasoning is gradually integrated into the prompts, the agents evolve from mechanical summarizers into analysts with genuine financial reasoning ability, capable of generating deeper insights and producing more accurate interpretations of market dynamics.
@@
4.2 Dual Trading Decision
In real-world trading, an account’s current holdings directly reflect its risk exposure. Professional traders adjust their buy–sell decisions dynamically based on existing positions to maintain a relative balance between risk and return. FinPos adopts a dual-agent framework that embeds position awareness into both the Direction Decision Agent and the Quantity and Risk Decision Agent. By integrating multi-source information retrieved from the memory layer, the dual-agent framework enables a decision-making mechanism that more closely mirrors practical trading logic.
4.2.1 The Direction Decision Agent
Incorporating position awareness requires integrating longer-term information. The direction decision agent combines structured memory (news, financial reports, and past actions) with the current portfolio state and explicitly specifies the strategic intent of each action—whether it represents a long-term position-building move or a short-term tactical adjustment to exploit local trends. This explicit articulation not only enhances interpretability for human supervisors but also provides downstream quantity agents with a clear strategic context. To prevent the agent from over-relying on short-term fluctuations, we design prompts that encourage it to factor in longer horizons, thereby gradually cultivating its long-term planning capability.
4.2.2 Quantity and Risk Decision Agent
FinPos introduces the Quantity and Risk Decision Agent, which determines the specific trade size after a directional decision has been made. This agent determines trade sizes through an explicit risk-aware mechanism that jointly incorporates the current position state, structured memory, and CVaR-based risk references (Conditional Value at Risk Rockafellar et al. (2000); see Appendix B.2), with per-step transaction sizes capped by the 95% CVaR to control exposure under volatility. This design ensures that every trade size is grounded in the agent’s current position and risk tolerance, making exposure control an explicit, first-class component of FinPos’s decision process.
4.3 Multi-Timescale Reward Design